OpenStudy (anonymous):

Cameron is replacing the roof on a gazebo. The roof is in the shape of a cone with an open bottom. The diameter is 14 feet and the slant height is 10 feet. What is the lateral area of the roof? Use 3.14 for pi.

OpenStudy (anonymous):

Please help...

jimthompson5910 (jim_thompson5910):

use the formula LSA = pi*r*s where r = radius s = slant height

OpenStudy (anonymous):

what's LSA

jimthompson5910 (jim_thompson5910):

LSA = lateral surface area

jimthompson5910 (jim_thompson5910):

in this case r = 7 (radius = half of the diameter) s = 10

OpenStudy (anonymous):

so my equation would look like this: LSA = 3.14 (7)(10)

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

evaluate that out

OpenStudy (anonymous):

219.8 is that right?

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

that helped alot

jimthompson5910 (jim_thompson5910):

you're welcome